mary works due north from home.her husband ben works due east.they leave from work at the same time. by the time mary is 5 miles from home the distance between them is one mile more than ben distance from home. how far is ben? __ miles ??
Answer by mananth(16949)

let Ben be x miles from home
then they will be x+1 miles away when Mary is 5 miles from home.
Since they are travelling east and north they form right triangle
Pythagoras theorem
..
(x+1)^2=x^2+5^2
x^2+2x+1=x^2+25
2x=24
x= 12 miles
ben is 12 miles from home
...
m.ananth@hotmail.ca